RC Meeting, Friday, October 01 , 2021

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meetings are at 10 am, remote

  • Bluejeans remote connection information:
 To join the Meeting:
 https://bluejeans.com/861439920
 To join via Room System:
 Video Conferencing System: bjn.vc -or-199.48.152.152
 Meeting ID : 861439920
 To join via phone :
 1)  Dial:
   +1.408.740.7256 (US (San Jose))
   +1.888.240.2560 (US Toll Free))
   +1.408.317.9253 (US (Primary, San Jose))
   (see all numbers - http://bluejeans.com/numbers)
 2)  Enter Conference ID : 861439920

Useful Links

Accelerator Status

  1. Accelerator started to deliver beam to Hall C ~5 pm yesterday and beam stability has been improving since (note by the end of Owl, accelerator has delivered mostly 60 uA, until recently, today morning, where they ramped up to 65 uA)

Hall C Status

  1. SHMS Q3 Interlock
    • At 7:27 this morning, SHMS Q3 Interlocked and did not reset. Amy Comer was contacted and interlock was cleared at 7:40 (thanks Amy !). As of 8:08, SHMS Q3 has been fully recovered and ready for resuming production --> SHMS Q3 Interlock

Current Run Plan

  1. Sep 30 Swing --> Oct 01 Owl/Day:
    • Contine data-taking at current PION-LT kinematic setting ( low eps, Q2=6 GeV2, W=3.14, x=0.39 )
      • Data collected as of 08:00 today: 73% of the total PAC approved beam charge accumulated. We still need to reach 85 % of beam charge before lowering the beam current to ~45 uA for the remainder of this kinematic until Saturday Afternoon.
      • From Jacob's estimates during Oct 01 Owl shift, we accumulated ~ 5 % / shift @ 60 uA of very stable beam. This would mean that we would reach ~85 % of beam charge by Saturday Oct 02 mid-Owl shift assuming beam keeps steady.
      • After 85% of beam charge has been accumulated ( this is contingent upon beam stability throughout weekend ), request to lower beam current to 45 uA to take the remaining data until TDB Day Shift --> https://logbooks.jlab.org/entry/3914672 ( PENDING )
  2. SHMS Rotation to 5.5 deg is contingent upon beam stability to Hall C over weekend (tentatively, Sunday Morning at 9:00 am )
    • Will contact Walter Kellner after RC meeting today to determine availability of techs on weekend to co-ordinate entrance to Hall. Will also need to contact Radcon today to give a specific time range at which rotation will take place on Sunday Morning. ( PENDING )
      • Rotation will require 2 beamline spotters (1 on top of HMS platform and one on the floor in-between HMS/SHMS)
      • Garth will volunteer to be a beamline spotter in case in case one of the techs cannot come in on a weekend, but an expert spotter (tech) is still required at the pivot
      • If possible, do rotation in "Controlled Access" (spotter in pivot will have to use ALARA to minimize time near radiation area), otherwise, will go into "Restricted Access" (takes a bit longer)

Current Issues

  1. Target ladder camera was identified to be bad: Joe Beaufait will purchase a replacement in the next funding cycle ( PENDING )
    • Still need to follow-up with Joe about when the purchase can be done.

Special Requests

  1. Contact Ellen Becker if there is an opportunity to access the hall (longer than 4 hours). This is for diagnosing the SHMS power supply.